# Coupling quantum-like cognition with the neuronal networks within generalized probability theory

### v1

@@ Abstract
The recent years are characterized by intensive applications of the methodology and mathematical apparatus of quantum theory, quantum-like modeling, in cognition, psychology, and decision making. In spite of the successful applications of this approach to a variety of psychological effects, e.g., the order, conjunction, disjunction, and response replicability effects, one may (but need not) feel dissatisfaction due to the absence of clear coupling to the neurophysiological processes in the brain. For the moment, this is just a phenomenological approach. In this paper we construct the quantum-like representation of the networks of communicating neurons. It is based not on standard quantum theory, but on generalized probability theory (GPT) with the emphasis of the operational measurement approach. We employ GPT’s version which is based on ordered linear state space (instead of complex Hilbert space). A network of communicating neurons is described as a weighted ordered graph that in turn is encoded by its weight matrix. The state space of weight matrices is embedded in GPT with effect-observables and state updates within measurement instruments theory. The latter plays the crucial role. This GPT based model shows the basic quantum-like effects, as e.g. the order, non-repeatability, and disjunction effects; the latter is also known as interference of decisions. This GPT coupling also supports quantum-like modeling in medical diagnostic for neurological diseases, as depression and epilepsy. Although the paper is concentrated on cognition and neuronal networks, the formalism and methodology can be straightforwardly applied to a variety of biological and social networks.
@@ 1 Introduction
The quantum information revolution (also known as the second quantum revolution) has not only the big technological impact (quantum computing, cryptography, simulators), but also the tremendous foundational consequences. One of them is understanding that the methodology and formalism of quantum theory can be applied outside of physics for modeling in cognition, psychology, decision making, biology including genetics and epigenetics, evolution theory, economics and finance, social and political sciences, game theory.111 See, for example, monographs [51, 20, 42, 9, 11, 43, 58, 12], the Palgrave handbook [44], and recent reviews [82, 57]. This area of research is known as quantum-like modeling (see, e.g., [81]).
It should be sharply distinguished from using of genuine quantum physics in cognition, as was done e.g. by Hameroff [40], Penrose [79], Vitiello [92]) and biology - “quantum biophysics” [6, 47]. The quantum-like models describe the behavior of macroscopic cognitive systems (e.g., human behavior) whose functioning can be modeled as information processing matching the laws of quantum information and probability and deviating from the classical laws.
@@
During the last 10–15 years quantum-like studies flowered by covering new domains of science, for example, genetics and biological evolution theory [7, 8, 14]. Hence, the second quantum revolution was completed by the quantum-like revolution [81]. The main challenge for the quantum-like approach is that the neurophysiological mechanism of generation of the macroscopic quantum-like behavior is still unknown. The experimental statistical data collected in cognitive psychology is characterized by, e.g., the order, conjunction, disjunction, and response replicability effects, and contexuality. Such data can be successfully described by the quantum formalism. [^2] See, for example, papers [36, 19, 41, 94, 95, 53, 61, 17, 75, 76, 77, 24, 1, 2, 16, 31, 32, 23, 13, 30, 46, 33]. However, we still don’t have a proper mathematical model coupling the neurophysiological processes in the brain with the quantum-like description. In particular, this problem was highlighted in article [21]. Nowadays this is not just a theoretical concern. Recently the quantum-like methodology started to be used for the medical diagnostics of neurological diseases [85, 88] with the first “quantum-like patent” [86]. Although this method gives excellent predictions, the absence of the neurophysiological basis slows its application in hospitals, since physicians want to see its neurophysiological justification.333 In principle, one might refer to quantum biophysics and to genuine quantum physical processing in the brain, e.g., in the microtubules [40, 79]. The quantum-like approach doesn’t contradict the quantum-brain hypothesis. But this hypothesis has so many drawbacks and it was heavily criticized by many authors, e.g., by Tegmark [90]. From our viewpoint, its main drawback is that in the quantum-brain theory the neurons are not the basic units of mental information processing. To find such units, one should go to finer scales of space and time, up to the Planck scale in Penrose’s coupling of consciousness with quantum gravity. However, we want to stay firmly on the neuronal processing basis of cognition and consciousness.
@@
In this paper we suggest a mathematical model of coupling between brain’s neurophysiology and the quantum-like project. It is based on the representation of functioning of brain’s neuronal networks within the generalized probability theory (GPT) (also known as the operational measurement/probability theory) [64, 38, 39, 25, 26, 67, 68, 72, 73], see [45, 80] for reviews. These networks are systems under analysis. Information exchange between neurons in a network $S$ is mathematically described by (generally directed) weighted graph $G_{S};$ neurons are mapped to the vertexes of $G_{S}$. We are well aware about difficulties of real experimental realization of such maps for the brain’s neuronal networks (see, e.g., [84]; so in our modeling we are still far from the experiment. Each graph $G_{S}$ corresponds to the information state of the network described by the weight matrix $\omega=(\omega_{ij})$. Matrix’s elements encode the strength of the signals between communicating neurons, e.g., the frequencies in the widely usedfrequency representation of the neural code [37] or correlations between the action The dynamics of weighted graphs mathematically describes the dynamics of network’s states. (But in this paper dynamics is not studied).
@@
This is the good place to mention some previous works on the neural network background of the quantum-like models. De Barros and Suppes [28] (see also [29]) pointed out that neural oscillators can produce interference patterns like that obtained from unitary evolution. Takahashi and Cheon [89] proposed a nonlinear neural network that used squared amplitudes to compute the probability of response in a quantum model of cognition. Busemeyer et al. [21] outlined one way that a neural based system could perform the computations required by applications of quantum probability to human behavior. Recently quantum-like modeling on treelike graphs started to be used in the medical research [85, 88] reflecting the hierarchy in information processing by the brain. Moreover, one of the first works in quantum-like modeling of cognition [49] was on quantum mechanics of the treelike graphs which are mathematically encoded by $p$-adic numbers. Later $p$-adics was applied to quantum-like modeling in research on petroleum reservoirs [63](with the treelike structure of their capillary networks). We also mention article [55] in that the quantum-like representation of mental information is seeded to the distribution of action potentials of neurons interacting within networks. The present paper was influenced by the discussion with Gregory Scholes on quantum-like states of the neural networks described by the special class of graphs [83, 5]. We consider arbitrary graphs.
@@
This paper is directed to those who are interested in the application of the quantum methods to cognition, psychology, and decision making. Generally researchers in this area are unaware about GPT. Moreover GPT is split in a few models that are apparently different but mathematically equivalent [67] (see [45, 80] for the reviews). We employ one special model; it is based on ordered linear spaces or so-called base norm spaces more precisely. To make the paper readable for experts in mathematical psychology, we briefly present the basics of GPT within the ordered linear space formalism in section 4.1.444It seems that the GPT approach can be more attractive for psychologists, than the standard formalism borrowed from quantum physics. GPT is “cleaned” from quantum mysteries and prejudices. This is simply a general mathematical formalism describing probabilities obtained in all kinds of observations done on all kinds of systems, physical, cognitive, social, financial. Before the formal introduction to GPT, we start by operating freely and as we hope clearly with the notions of GPT, and in section 3 we present the model bridging quantum-like cognition and brain’s networks of communicating neurons. Then we come back to this model in section 5 to show the order, interference (disjunction), and non-repeatability effects. For reader’s convenience the appendix starts with the GPT description of the conventional quantum probability (in complex Hilbert space).
@@
This paper is conceptual, its aim is to present the concept of GPT coupling of neuronal activity with the quantum cognition project. In further papers we shall go deeper into the technical details. Although the paper is concentrated on cognition and neuronal networks, the formalism and methodology can be straightforwardly applied to a variety of biological and social networks with coupling to quantum-like modeling in genetics and social science [59, 4].
@@ 2 Graph representation of the networks of communicating neurons
The voltage-dependent channels endow the neuron with active signaling property. In particular, they enable a neuron to generate nerve impulses, also know as action potentials. In spite of variation in duration and amplitude as well as shape, they are typically treated as identical events. The duration of an action potential (approximately 1 ms) is ignored, an action potential sequence, or spike train, is represented by a series of all-or-none point events.
We note that each neuron has, on average, about 7,000 synaptic connections with other neurons. Axons vary in length, some of them are a millimeter or so, but the axons which go from the brain down the spinal cord, can extend for more than a meter. An axon has numerous side branches called axon collaterals, that is one neuron can send information to several others. The collaterals, as the roots of a tree, split into smaller extensions called terminal branches.
Consider some network $S$ in the brain consisting of $N$ neurons, e.g., a network performing some brain function or even the network of all neurons in the brain. Neuronal networks are systems of our quantum-like model. The symbol $S$ is used to denote a neuronal network of communicating neurons.
In the graph representation of $S$, neurons are mapped to the graph’s vertexes, connections between neurons are mapped to graph’s edges. We treat connections from the information viewpoint, so edges need not be physical edges consisting of axons and synapses. For example, edges (connections) between pairs of neurons are established by correlations between their signaling. We recall that if no two edges have the same endpoints we say there are no multiple edges, and if no edge has a single vertex as both endpoints we say there are no loops [91]. A graph $G_{S}$ representing network $S$ can have multiple edges and loops.
A directed graph consists of a set of vertices and a set of directed edges that are ordered pairs of vertices. A directed graph allows loops and allows two edges joining the same vertex, but going in the opposite direction. More than one edge going in the same direction between vertices, however, is not allowed [91]. The main model considered in this article (Model 1, section 3) is based on the directed neuronal graphs. Other two models are based on undirected graphs with loops, but without multiple edges.
@@ 2.1 Model 1: directed graph for neurons signaling
Here $0\leq w_{ij}\leq 1$. We note that generally $w_{ii}>0$ that is a neuron can send signals to itself. The GPT representation for this neuronal model will be presented in section 3.
@@ 4 Generalized probability theory (operational measurement theory)
GPT was created in 1970th as an attempt to find the roots of the quantum formalism. Within GPT quantum formalism based on the complex Hilbert space is just a special version of the general operational formalism for all kinds of measurements. It unifies the classical and quantum measurement theories and opens the door to the variety of quantum-like formalisms. It is clear that probabilistically non-classical properties of mental phenomena need not imply that they should be described by the quantum (complex Hilbert space) formalism, quantum probability, information and measurement theory. Some GPT-based models can match better cognitive information processing and decision making; this was highlighted in [60]. [^5] In 2015 one of the coauthors (AKH) by staying in Vienna had long discussions with Anton Zelinger on possible applications of quantum theory to cognition. These discussion were initiated by conversations about his debates with Dalai Lama on spiritual versus quantum mechanical phenomena. He was strongly in favor of appealing to generalized theories which are neither classical nor quantum; see also his preface to book [81]. He also guessed that it may be that final quantum-like theory of cognition would be mathematically so complicated and exotic that the complex Hilbert formalism would not more look as complex and advanced. But this is not the case for the model suggested in this article which is “less quantum” than the quantum mechanics.
In this paper we employ GPT based on theory of ordered linear spaces [25, 26, 67, 68, 72, 73, 74]. In contrast to the majority of GPT-studies, we emphasize not the structures of the state and effect spaces, but the calculus of measurement instruments giving the probabilities (via generalized Born’s rule) and states’ updates generated by the measurements feedback. This approach to GPT started by Davies and Lewis [25] and known as quantum instruments theory, and it has been developed by Ozawa [67, 68, 69, 70, 71, 72, 73, 74], see also the recent article by Khrennikov [60] in that this approach was restructured in the contextual framework. Such contextual viewpoint of GPT and quantum instruments theory enlightens the possibility of their applications to cognitive psychology, decision making, social science [60]. GPT research which is concentrated on generalization of the notions of state and observable was started in [64] and essentially developed by Gudder [38, 39]. Ozawa [67]. established the equivalence between Gudder’s and Davies–Lewis’ approaches.
@@ 4.3 Instruments
As was emphasized in [76], the basis of any measurement model is its instrument calculus. One can’t start from observables, as often done, in quantum information theory, where one is typically fine by introducing just observables. An observable is only a byproduct, or an “effect”, of an instrument, and the statistics of a measurement is fully described by the corresponding instrument.
@@ 4.5 Order effect
The question order effect is well studied in psychology and sociology, e.g., for social opinion pools [65]. In classical probability theory the order effect is absent. In [94, 95] it was represented in the quantum-like framework based on the projection type (von Neumann [93]) state updates. But then in [53] it was shown that such projection description can’t be combined with another psychological effect, the response replicability effect. To resolve this problem, in [75, 76] the calculus of quantum instruments was employed. We now consider the order effect within GPT.

### v2

@@ Abstract
The past few years have seen a surge in the application of quantum theory methodologies and quantum-like modeling in fields such as cognition, psychology, and decision-making. Despite the success of this approach in explaining various psychological phenomena — such as order, conjunction, disjunction, and response replicability effects — there remains a potential dissatisfaction due to its lack of clear connection to neurophysiological processes in the brain. Currently, it remains a phenomenological approach. In this paper, we develop a quantum-like representation of networks of communicating neurons. This representation is not based on standard quantum theory but on generalized probability theory (GPT), with a focus on the operational measurement framework. Specifically, we use a version of GPT that relies on ordered linear state spaces rather than the traditional complex Hilbert spaces. A network of communicating neurons is modeled as a weighted directed graph, which is encoded by its weight matrix. The state space of these weight matrices is embedded within the GPT framework, incorporating effect-observables and state updates within the theory of measurement instruments — a critical aspect of this model. This GPT-based approach successfully reproduces key quantum-like effects, such as order, non-repeatability, and disjunction effects (commonly associated with decision interference). Moreover, this framework supports quantum-like modeling in medical diagnostics for neurological conditions such as depression and epilepsy. While this paper focuses primarily on cognition and neuronal networks, the proposed formalism and methodology can be directly applied to a wide range of biological and social networks.
@@ 1 Introduction
The quantum information revolution (commonly referred to as the second quantum revolution) has had a profound impact, not only technologically — ushering in advancements in quantum computing, cryptography, and simulation — but also in foundational understanding. One key revelation is that the methodology and formalism of quantum theory can be extended beyond physics to model phenomena in diverse fields, such as cognition, psychology, decision-making, biology (including genetics and epigenetics), evolutionary theory, economics, finance, social and political sciences, and game theory. This interdisciplinary area of research is broadly referred to as quantum-like modeling (e.g., [81]). [^1] For examples, see monographs [51, 20, 42, 9, 11, 43, 58, 12], the Palgrave Handbook [44], and recent reviews [82, 57]. It should be clearly distinguished from the use of genuine quantum physics in cognition, as explored by researchers such as Hameroff [40], Penrose [79], and Vitiello [93], as well as in biology, particularly in the framework of quantum biophysics [6, 47]. Quantum-like models, on the other hand, describe the behavior of macroscopic cognitive systems (e.g., human behavior), where the functioning of these systems can be modeled as information processing that adheres to the laws of quantum information theory and probability, deviating from classical laws.
@@
In the past 10–15 years, quantum-like studies have expanded into new domains of science, such as genetics and biological evolution theory [7, 8, 14]. As a result, the second quantum revolution has been complemented by the quantum-like revolution [81]. A key challenge for the quantum-like approach lies in the fact that the neurophysiological mechanisms responsible for generating macroscopic quantum-like behavior remain unknown. The experimental statistical data collected in cognitive psychology exhibits phenomena like the order, conjunction and disjunction, response replicability effects, and contextuality. Such data can be successfully described using quantum formalism.  [^2] See, for example, the works cited in [36, 19, 41, 95, 96, 53, 61, 17, 75, 76, 77, 24, 1, 2, 16, 31, 32, 23, 13, 30, 46, 33, 91]. However, there is still no proper mathematical model that links the neurophysiological processes in the brain with a quantum-like description. This issue was highlighted in [21]. Currently, this is not merely a theoretical concern. Quantum-like methodology has recently been applied in the medical diagnosis of neurological diseases [85, 88], and the first “quantum-like patent” has been granted [86]. While this method provides highly accurate predictions, the lack of a neurophysiological foundation hinders its widespread use in hospitals, as physicians require a neurophysiological justification for its application.  [^3] One might refer to quantum biophysics or to genuine quantum physical processing in the brain, such as in the microtubules [40, 79]. And the quantum-like approach does not contradict the quantum-brain hypothesis, despite the latter’s numerous drawbacks, as pointed out by many critics, including Tegmark [90]. From our perspective, the main issue with the quantum-brain theory is that it does not consider neurons as the fundamental units of mental information processing. To identify such units, one would need to explore finer scales of space and time, possibly down to the Planck scale, as in Penrose’s theory of consciousness and quantum gravity. However, we aim to focus on neuronal processing as the core basis for cognition and consciousness.
@@
In this paper, we propose a mathematical model for the coupling between the brain’s neurophysiology and the quantum-like framework. This model is based on the representation of brain neuronal network functioning within generalized probability theory (GPT) (also known as operational measurement/probability theory) [64, 38, 39, 25, 26, 67, 68, 72, 73]; for reviews, see [45, 80]. These neuronal networks are considered as systems under analysis. Information exchange between neurons in a network is mathematically described by a (generally directed) weighted graph $G_{S},$ where the neurons are mapped to the vertices of $G_{S}.$ We acknowledge the challenges in experimentally realizing such mappings for the brain’s neuronal networks (see, e.g., [84]), and thus, our modeling remains far from experimental implementation. Each graph $G_{S}$ corresponds to the information state of the network, which is described by the weight matrix $\omega=(\omega_{ij}).$ The elements of this matrix encode the strength of signals between communicating neurons, such as the frequencies used in the widely adopted frequency representation of the neural code [37] or correlations between neuronal activations. The dynamics of the weighted graph mathematically describes the evolution of the network’s states. (Note that the dynamics are not studied in this paper.)
@@
This is an opportune moment to highlight prior research on the neural network foundations of quantum-like models. De Barros and Suppes [28] (see also [29]) demonstrated that neural oscillators can generate interference patterns analogous to those observed in unitary evolution. Takahashi and Cheon [89] proposed a nonlinear neural network model that uses squared amplitudes to compute response probabilities within a quantum cognitive framework. Busemeyer et al. [21] explored how a neural-based system could perform the computations required for applying quantum probability to human behavior. More recently, quantum-like modeling on treelike graphs has gained traction in medical research [85, 88], reflecting the hierarchical nature of information processing in the brain. Notably, one of the earliest works in quantum-like modeling of cognition [49] focused on the quantum mechanics of treelike graphs, which are mathematically represented using $p$-adic numbers. This approach was later extended to quantum-like modeling in studies of petroleum reservoirs [63], where the treelike structure of capillary networks was analyzed. Additionally, in [55], mental information was represented quantum-like through the distribution of action potentials in neuron networks. The present work also draws inspiration from discussions with Gregory Scholes on quantum-like states in neural networks represented by specialized classes of graphs [83, 5]. However, our approach considers arbitrary graphs.
@@
This paper is intended for those interested in applying quantum methods to cognition, psychology, and decision-making. Researchers in these areas are generally unaware of GPT. Moreover, GPT is divided into several models that may appear to be different, but are mathematically equivalent [67] (for reviews, see [45, 80]). In this paper, we employ a specific model based on ordered linear spaces, or more precisely, base norm spaces. To make the paper accessible to experts in mathematical psychology, we briefly introduce the basics of GPT within the ordered linear space formalism in section 4.1.  [^4] It seems that the GPT approach may be more appealing to psychologists than the standard formalism borrowed from quantum physics. GPT is “cleansed” of the quantum mysteries and prejudices, offering a more straightforward general mathematical framework that describes probabilities in all kinds of observations—whether physical, cognitive, social, or financial.
Before formally introducing GPT, we begin by intuitively exploring and, we hope, clearly using the concepts of GPT. In section 3, we present a model that bridges quantum-like cognition and the brain’s networks of communicating neurons. We return to this model in section 5 to demonstrate the effects of order, interference (disjunction), and non-repeatability.
@@
This paper is conceptual in nature, aiming to introduce the concept of coupling neuronal activity with the quantum cognition project through GPT. Further papers will delve deeper into the technical details. While this paper focuses on cognition and neuronal networks, the formalism and methodology presented here can be easily applied to a variety of biological and social networks, including quantum-like modeling in genetics, medicine, and social science [59, 85, 88, 4].
@@ 2 Graph representation of the networks of communicating neurons
Voltage-dependent channels endow neurons with active signaling capabilities. Specifically, they enable neurons to generate nerve impulses, also known as action potentials. Despite variations in duration, amplitude, and shape, action potentials are typically treated as identical events. The duration of an action potential (approximately 1 ms) is often ignored, and an action potential sequence, or spike train, is represented as a series of all-or-none point events.
It is important to note that each neuron, on average, has about 7,000 synaptic connections with other neurons. Axons vary in length; some are only a millimeter long, while those extending from the brain down the spinal cord can be over a meter in length. An axon has numerous side branches, known as axon collaterals, allowing one neuron to send information to several others. These collaterals, much like the roots of a tree, split into smaller extensions called terminal branches.
Now, consider a network $S$ in the brain consisting of $N$ neurons, such as a network performing a specific brain function or even the entire network of neurons in the brain. Neuronal networks are the systems we model in our quantum-like framework. Throughout the paper the symbol $S$ denotes a neuronal network of communicating neurons.
In the graph representation of $S,$ neurons are mapped to the vertices of the graph, and the connections between neurons are represented as the graph’s edges. From an information-theoretic perspective, these edges need not correspond to physical connections, such as axons and synapses. For example, edges (connections) between pairs of neurons may be established based on correlations in their signaling.
We recall that if no two edges share the same pair of endpoints, the graph is said to have no multiple edges. Additionally, if no edge connects a vertex to itself (i.e., no edge has the same vertex as both endpoints), the graph has no loops [92]. A graph $G_{S}$ representing network $S$ may contain multiple edges and loops.
A directed graph consists of a set of vertices and a set of directed edges, which are ordered pairs of vertices. A directed graph allows loops and permits two edges joining the same vertex but going in opposite directions. However, multiple edges going in the same direction between vertices are not allowed [92]. The primary model discussed in this article (Model 1, section 3) is based on directed neuronal graphs. The other two models are based on undirected graphs with loops, but without multiple edges.
@@ 2.1 Model 1: directed graph for neurons signaling
Here $0\leq w_{ij}\leq 1$. We note that generally $w_{ii}>0$ that is a neuron can send signals to itself. The GPT representation for this neuronal model will be presented in section 3. See Fig. 1 for illustration.
0,1; Refer to caption
Figure 1: An abstract representation of two neurons network. A directed graph representing neuron signaling implies that generally $W_{12}\neq W_{21}.$ Each neuron has two axon collaterals: one ending in a synapse on a dendrite of the other neuron, and the second collateral connecting back to a dendrite of the same neuron. The insert depicts the synaptic cleft.
@@ 4 Generalized probability theory (operational measurement theory)
GPT was developed in the 1970s as an attempt to understand the foundations of quantum formalism. Within the framework of GPT, the quantum formalism based on complex Hilbert spaces is viewed as a special case of the general operational formalism for all types of measurements. GPT unifies classical and quantum measurement theories and paves the way for a variety of quantum-like formalisms.
It is important to note that the probabilistically non-classical properties of mental phenomena do not necessarily imply that they should be described using quantum formalism (such as the complex Hilbert space), or quantum probability, information, and measurement theory. Some GPT-based models may better align with cognitive information processing and decision-making, as highlighted in [60].  [^5] In 2015, one of the co-authors (AKH) had extensive discussions in Vienna with Anton Zeilinger about possible applications of quantum theory to cognition. These discussions were prompted by Zeilinger’s debates with the Dalai Lama on the relationship between spiritual and quantum mechanical phenomena. Zeilinger strongly advocated for generalized theories that are neither purely classical nor quantum. He also suggested that the ultimate quantum-like theory of cognition might be so mathematically complex and exotic that the complex Hilbert formalism would no longer appear as advanced. However, this is not the case for the model proposed in this article, which is “less quantum” (“more classical”) than traditional quantum mechanics.
In this paper, we employ GPT based on the theory of ordered linear spaces [25, 26, 67, 68, 72, 73, 74]. Unlike the majority of GPT studies, we focus not on the structures of state and effect spaces, but on the calculus of measurement instruments that provide probabilities (via the generalized Born’s rule) and the updates to states generated by the feedback from measurements. This approach, introduced by Davies and Lewis [25] and known as quantum instruments theory, has been further developed by Ozawa [67, 68, 69, 70, 71, 72, 73, 74]. A recent article [60] restructured this approach within a contextual framework. The contextual perspective of GPT and quantum instruments theory illuminates their potential applications to cognitive psychology, decision-making, and social science [60].
GPT research, which focuses on generalizing the concepts of state and observable, was initiated in [64] and significantly developed by Gudder [38, 39]. Ozawa [67] established the equivalence between Gudder’s and Davies–Lewis’ approaches.
@@ 4.3 Instruments
As emphasized in [76], the foundation of any measurement model lies in its instrument calculus. One cannot start with observables, as is often done in quantum information theory, where the introduction of observables is typically sufficient. In reality, an observable is merely a byproduct, or an “effect,” of an instrument, and the statistics of a measurement are fully described by the corresponding instrument.
@@ 4.5 Order effect
The question order effect is well-studied in psychology and sociology, particularly in the context of social opinion polls [65]. In classical probability theory, the order effect is absent. In [95, 96], this effect was represented within a quantum-like framework based on projection-type (von Neumann [94]) state updates. However, in [53], it was demonstrated that such a projection-based description cannot be combined with another psychological effect, the response replicability effect. To address this issue, the calculus of quantum instruments was employed in [75, 76]. We now examine the order effect within the framework of GPT.